Circular No. 6970

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998cx IN NGC 6209 A. Wassilieff, Palmerston North, New Zealand, reports his discovery

of an apparent supernova (mag ∼ 17.8) on July 4.412 UT with a 0.36-m f/4.3 reflector (+ unfiltered SBIG ST-7 CCD). G. Nankivell, Wainuiomata, Lower Hutt, N.Z., provides the following measurement for the position of SN 1998cx from a July 4 image by Wassilieff: α = 16^{h}54^{m}55^{s}.31, δ = −72^{o}35′24′′.0 (equinox 2000.0), which is 10′′.9 west and 12′′.1 south of the nucleus of NGC 6209. Further rough magnitudes from Wassilieff: July 5, 17.8; 16, 18.3. SN 1998cx does not show on a CCD image taken on 1993 Aug. 18, nor is it evident on a second-generation Digital Sky Survey image.

SUPERNOVA 1998cu IN IC 1525 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Center for Astro-

physics, report that a spectrum of SN 1998cu (cf. IAUC 6958) taken by P. Berlind at the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph) on July 15.4 UT exhibits a prominent P-Cyg Hα feature, along with Hβ and Fe II absorption, indicating that the supernova is of type II after maximum light. The photospheric expansion velocity measured from the Hα profile is 5200 km/s, based on the NASA/IPAC Extragalactic Database recession velocity for IC 1525 of 5022 km/s.

CH CYGNI M. Karovska, Harvard-Smithsonian Center for Astrophysics; C. Carilli,

National Radio Astronomy Observatory; and J. Mattei, American Association of Variable Star Observers, report that they have been monitoring the radio brightness changes of the symbiotic system CH Cyg at 8, 15 and 22 GHz since Oct. 1996, following an extreme fading of the visual brightness, to mag ∼ 10.4, a few months earlier (IAUC 6426): “The radio brightness remained relatively constant at ∼ 6–8 mJy at all frequencies until Apr.– May 1997, when we detected a sudden flux increase by ∼ 50 percent at several frequencies; the visual magnitude was then ∼ 9.4. In addition to the flux measurements, we obtained high-angular-resolution maps of CH Cyg at several frequencies showing that the source is resolved. Recent radio maps, obtained from observations carried out during May-June 1998, show a significant north-south elongation (of ∼ 0′′.3), possibly indicating a jet formation.”

1998 July 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green